The dimensions of a rectangular flower garden were 8 m by 15 m . Each dimension was increased by the same amount. The garden then had an area of 198 m2. Find the dimensions of the new garden.
Answer by tquon(6)

If each side was raised by the same amount, x, then (8+x)*(15+x)=198. So 120+23x+x^2=198. subtract 198 from each side to get x^2+23x-78. You can factor this to get (x-3)*(x+26). Your possible solutions for x are -26 and 3, but since it cant be increased by a negative number, 3=x. so back to the original equation substituting in for x, (8+3)*(15+3)=11*18=198. so the new dimensions are 11mx18m.